Let R be a commutative Ring with identity.let I and J be ideals of R such that I+J=Show that IJ= intersection of I and J.
1
Expert's answer
2020-10-20T17:50:39-0400

Let "R" be a commutative ring with identity and "I" and "J" be ideals of "R" such that "I+J=R". Let us prove that "IJ=I\cap J".


Since "I" is an ideal, "IJ\subset IR\subset I". Since "J" is an ideal, "IJ\subset RJ\subset J". Therefore, "IJ\subset I\cap J".


On the other hand, taking into accounr that "R" is a commutative ring with identity, we conclude that


"I\cap J\subset (I\cap J)R\subset (I\cap J)(I+J)= (I\cap J)I+ (I\cap J)J\subset JI+IJ=IJ+IJ=IJ".


Consequently, "IJ=I\cap J".